### v2

# Slot-BERT: Self-supervised Object Discovery in Surgical Video

Abstract
Object-centric slot attention is a powerful framework for unsupervised learning of structured and explainable representations that can support reasoning about objects and actions, including in surgical video. While conventional object-centric methods for videos leverage recurrent processing to achieve efficiency, they often struggle with maintaining long-range temporal coherence required for long videos in surgical applications. On the other hand, fully parallel processing of entire videos enhances temporal consistency but introduces significant computational overhead, making it impractical for implementation on hardware in medical facilities. We present Slot-BERT, a bidirectional long-range model that learns object-centric representations in a latent space while ensuring robust temporal coherence. Slot-BERT scales object discovery seamlessly to long videos of unconstrained lengths. A novel slot contrastive loss further reduces redundancy and improves the representation disentanglement by enhancing slot orthogonality. We evaluate Slot-BERT on real-world surgical video datasets from abdominal, cholecystectomy, and thoracic procedures. Our method surpasses state-of-the-art object-centric approaches under unsupervised training achieving superior performance across diverse domains. We also demonstrate efficient zero-shot domain adaptation to data from diverse surgical specialties and databases.
1 Introduction
Research suggests that humans learn to perceive the world through object-specific grouping and association (Kahneman et al., 1992; Tenenbaum et al., 2011). This ability constructs representations akin to object files that bind and track features of an object over time thereby enabling more efficient use by downstream cognitive tasks. Inspired by this cognitive process, unsupervised object-centric learning aims to learn explainable and adaptable representations from unlabeled datasets (Burgess et al., 2019; Greff et al., 2019; Locatello et al., 2020; Greff et al., 2020).
While object-centric learning comes in different forms, one particularly effective apporach incorporates the inductive bias of grouping or binding low-level, unstructured perceptual activations into a set of vectors known as slots. Each slot encapsulates a higher-level compositional entity, such as an object. This grouping or binding process emerges from architectural priors combined with self-supervised learning techniques, including attention and auto-encoding through specialized encoder-decoder architectures. Substantial progress has been made in this field for both image (Seitzer et al., 2023; Fan et al., 2024; Mansouri et al., 2023; Jiang et al., 2023; Wu et al., 2023b) and video (Weis et al., 2020; Kipf et al., 2022; Aydemir et al., 2023; Zadaianchuk et al., 2023; Wu et al., 2023a; Biza et al., 2023; Lee et al., 2024; Qian et al., 2023; Bao et al., 2022b, 2023; Singh et al., 2022, 2024) processing.
In object-centric image processing, features like color and semantic embeddings from pre-trained models are used to learn pixel-to-object assignments. Methods proposed for object-centric processing of video add additional cues about the temporal consistency of objects, including optical flow (Kipf et al., 2022) and between-frame similarities (Zadaianchuk et al., 2023) or depth maps (Elsayed et al., 2022). While beneficial, these additional cues can lead to increased computational complexity and a higher risk of error accumulation. For instance, optical flow often fails with static or deformable objects and large inter-frame displacements, while depth maps may be not available for certain video domains and unreliable in low-light or low-contrast settings. Feature similarity can address certain challenges but may result in homogeneity when training on smaller datasets as the similarity map is agnostic to the semantics of these features.
These issues become especially pronounced in surgical videos, where temporal dynamics can be highly complex. Instruments and tissues in surgical videos often move at different speeds, and the visibility of specific objects can fluctuate significantly over the course of a procedure. Moreover, many current object-centric learning approaches for videos fail to fully exploit the sequential structure of datasets with longer videos, primarily due to their reliance on RNN-based models (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022) that struggle with learning from long-range temporal coherence. Employing slot-attention mechanisms to process entire video sequences, as proposed in (Singh et al., 2024), could help overcome some challenges but may struggle with scalability for longer videos in real-world datasets due to computational limitations.
In this paper we propose an architecture that is both easy to train and excels in long-range bidirectional temporal reasoning to efficiently handle longer video sequences. We introduce (Slot-BERT), an unsupervised object-centric slot attention model that performs temporal reasoning over slots using a bidirectional transformer encoder. The model treats slots learned from video frames as foundational visual concepts, similar to how text embeddings represent words. By adapting the BERT transformer encoder from natural language processing to process encoded video slots (akin to a sequence of word embeddings), Slot-BERT learns to reconstruct video feature maps, providing implicit supervision for object discovery through masked auto-encoding of slots. This approach implements a temporal bidirectional self-attention mechanism, enabling effective video representation learning. Furthermore, by learning to reconstruct masked slots, Slot-BERT can predict future slots, and we can use this functionality for future initialization of slots that improves the accuracy when applied on longer sequences. We also propose a video slot contrastive loss that increases the independence between slots within a video by maximizing orthogonality in the latent vector space. This regularization leads to a more distinct representation of concepts and increases the precision of segmentation maps.
To summarize, our contributions are as follows: (1) We introduce Slot-BERT, a novel object-centric self-supervised representation learning model based on bidirectional temporal reasoning across video frames using a transformer architecture. (2) We introduce slot-contrastive loss, specifically designed for slot attention, to improve orthogonality between slots. (3) We demonstrate superior temporal coherence, bidirectional reasoning, and zero-shot generalization compared to state-of-the-art methods across four surgical video datasets from three different domains: abdominal, cholecystectomy, and thoracic surgery. (4) Our model addresses the limitations of fully parallelized architectures, scales effectively to longer videos, and can run on affordable hardware.
2 Related Work
2.1 Self-supervised object-centric learning
Learning of object-centric representations can demonstrably improve the sample efficiency and generalization of vision and dynamics models in compositional domains. In order to discover objects from images or videos without supervision, the learning architecture must bind the distinct features that belong to particular objects and create representations with activation patterns that are distinct for object instances. Various approaches have been proposed to achieve this goal. Early works focused on variational auto-encoders (VAEs) (Kingma & Welling, 2014) and disentangled representation learning (Mathieu et al., 2019; Eastwood & Williams, 2018; Kim & Mnih, 2018; Higgins et al., 2017). Other efforts employed iterative attention mechanisms, such as Capsule Networks (Sabour et al., 2017; Hinton et al., 2018; Tsai et al., 2020), contrastive learning approaches (Kipf et al., 2020; Hénaff et al., 2022; Xu et al., 2022), or complex-valued auto-encoders that implicitly encode temporal correlation (Löwe et al., 2022). Beyond learning generalized representations, reconstructive and generative object-centric models also show the ability to segment objects in scenes (Lin et al., 2020; Van Steenkiste et al., 2020; Greff et al., 2019; Engelcke et al., 2020; Burgess et al., 2019).
Slot Attention (SA) (Locatello et al., 2020; Kori et al., 2024), a recently proposed iterative attention mechanism, proved effective in grouping latent representations of objects into a number of slots, and can scale to real-world scenes by reconstructing pre-trained features from foundation models like DINO (Caron et al., 2021) or MAE (He et al., 2022). SA can use multiple modalities to discover objects, such as optical flow (Kipf et al., 2022), depth maps (Elsayed et al., 2022), temporal feature similarity (Zadaianchuk et al., 2023), or 3D point clouds (Ibrahim et al., 2023). Furthermore, SA can embed object entities in an identifiable manner (Kori et al., 2024), with promising results demonstrated on synthetic data. It has also been applied to improve Vision-Language Modeling (Xu et al., 2024) and controllable synthetic image generation (Jiang et al., 2023; Wu et al., 2023b).
In video processing, association of slots across time requires initialization and SA on video typically uses RNN-like architectures (Kipf et al., 2022; Zadaianchuk et al., 2023; Singh et al., 2022; Elsayed et al., 2022) that support iterative initialization of slots. These approaches however suffer from instability when training on longer sequences and in scenes with a limited temporal coherence. A parallelized approach, such as (Singh et al., 2024), processes the entire video sequence as a batch, which comes at the cost of computational efficiency and makes it less suitable for real-world video applications. Our SA architecture builds on the RNN framework by incorporating a temporal fusion transformer to enhance long-range bidirectional reasoning. This novel design improves temporal coherence and maintains scalability for longer video sequences while addressing the limitations of prior methods.
2.2 Masked Information Encoding
Masked visual modeling has been proposed to learn effective visual representations, starting with training of denoising auto-encoders where masks were treated as noise (Vincent et al., 2010). Transformer architectures (Vaswani et al., 2017) brought significant progress in Natural Language Processing (NLP) (Devlin et al., 2019; Radford, 2018) and vision (Dosovitskiy et al., 2021; Arnab et al., 2021) thanks to learned dot product attention for long-range encoding of co-occurrences. Building upon the success of masked self-attention in GPT (Radford, 2018), iGPT learned to predict sequences of pixels (Chen et al., 2020a) or masked tokens as in the ViT model (Dosovitskiy et al., 2021)
The success of these vision transformers inspired a number of other transformer-based architectures for masked visual modeling (Bao et al., 2022a; Dong et al., 2023; Wei et al., 2022). BEiT (Bao et al., 2022a) and BEVT (Wang et al., 2022) followed the language model BERT (Devlin et al., 2019) learning visual representations from images and videos by predicting discrete tokens (Ramesh et al., 2021). Image MAE (He et al., 2022) and Video MAE (Tong et al., 2022) used asymmetric encoder-decoder architectures for masked image modeling based on plain ViT backbones. MaskFeat (Wei et al., 2022) proposed reconstructing HOG features of masked tokens to perform self-supervised pre-training in videos.
Masked-out features or tokens have also been applied to weakly supervised settings, where class labels were provided while the objective was to produce attention maps for detection and segmentation (Hou et al., 2018; Lee et al., 2019). Inspired by BERT, our method treats slots encoded from images as analogous to word embeddings, incorporating a transformer to process video slots. The training objective is to recover image features when slots are partially masked out. Following this strategy, the transformer module can perform reasoning across time instead of merely duplicating information from individual slots.
2.3 Unsupervised Object Detection and Segmentation in Video
Unsupervised object detection and segmentation in videos often rely on motion cues to identify objects and regions of interest. Motion-based approaches have been extensively applied to video segmentation (Fragkiadaki et al., 2015; Kossen et al., 2020; Ponimatkin et al., 2023; Liu et al., 2021; Karazija et al., 2022; Choudhury et al., 2022). For example, Fragkiadaki et al. (Fragkiadaki et al., 2015) ranked segment proposals by combining optical flow with static boundary information. Similarly, Tokmakov et al. (Tokmakov et al., 2017) utilized optical flow to capture motion cues; however, their method struggles to segment static objects due to the lack of detailed spatial information.
To address these challenges, MATNet (Zhou et al., 2020) incorporated motion information to enhance spatiotemporal object representation. While this fusion of static and motion data improved performance, it encountered difficulties with complex moving backgrounds and was highly dependent on the accuracy of optical flow maps. In the domain of surgical videos, Sestini et al. (Sestini et al., 2023) employed a teacher-student network to refine optical flow-based segmentation, improving accuracy in these specialized scenarios.
Beyond motion-based methods, Croitoru et al. (Croitoru et al., 2019) proposed using Principal Component Analysis (PCA) to isolate foreground objects in a fully unsupervised manner. For cases where the segmentation of the first frame is available, tracking-based methods (Uziel et al., 2023; Cheng et al., 2023) can effectively propagate segmentation across video frames. To effectively address the issue of moving background and uncertainty of object presence in surgical videos, an approach using feature ranking and knowledge distillation is proposed (Liao et al., 2024), however this approach needs weak supervision using video-level class labels.
Our work focuses on object-centric learning that produces explainable representations and slot-based segmentation masks at the same time. While we leverage motion cues for object discovery through inductive biases, our method can learn moving and static objects using a unified framework. By relying solely on a feature reconstruction objective it simplifies the training process while maintaining robustness in handling long-range temporal dependencies. This design allows for more accurate segmentation and object discovery across diverse video domains.
3 Methods
3.1 Overview
We base our object-centric representation learning framework (Figure 1) on slot attention with RNN (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022). A video sequence = , where is an image at time step of size with channels, and is a fixed length of the video clip that our model operates on, is encoded using a self-supervised feature extraction model to obtain . Following a recurrent iterative attention approach we obtain the slots representation = , where are the latent space slots that embed objectness information of image . A Temporal Slot Transformer (TST) module feeds to a masked transformer encoder which allows elements to attend to each other and aggregate the long-distance information to form . An image decoder then recurrently maps each element of to the video encoding space . The main training objective is to minimize the distance between the reconstructed feature and the original feature representation .
3.2 Temporal recurrence of object-centric slot encoder
We encode patches of size from individual input frames using the Vision Transformer (ViT) encoder, producing a stack of patch feature embeddings , where and .
Features representing one individual image are grouped by the slot attention module (Locatello et al., 2020) into spatial groupings through iterative competitive attention and encoded into slots . Let denote the key and value transformation matrices acting on , and the query transformation matrix acting on . The iterative slot update function is based on dot attention:
where is the query vector, is the key vector, is the value vector, and is the attention matrix.
Slot attention can be seen as a version of self-attention (Vaswani et al., 2017) constrained by an object-centric inductive bias that is much more computationally efficient as the attention query is much smaller ( , e.g. for patches, for objects in an image).
The assignment of slots across video frames is initialized by a RNN-like computation where slots = for -th frame are initialized by the previous frame’s final slot estimation. First frame’s slots are initialized randomly by sampling from a standard Gaussian. In this way, the permutation of groupings of objects across slots is likely to remain the same at neighboring frames. However, this assignment could not guarantee temporal consistency for longer video sequences. We thus introduce the TST module to enhance long-range bidirectional reasoning along the temporal dimension.
3.3 Temporal slot transformer
The TST module facilitates interactions across time steps and enables robust temporal consistency by leveraging positional embeddings and masked training. It processes a sequence of slots , where is the number of frames, is the number of slots, and is the slot dimension. The module comprises the following components:
Temporal Positional Embeddings
: A learnable positional embedding is added to the slots to encode temporal information:
Masked Transformer Encoder
: To promote robust learning, and allow TST to have bidirectional reasoning ability (i.e., can predict previous and next slots) we employ random masking during training. For a given masking ratio , a subset of frames is masked by setting their corresponding slot values to zero:
where is the binary mask, and represents element-wise multiplication. The masked slots are then processed through a Multi-head Transformer encoder:
Our masked transformer implementation follows previous work in language modeling (Devlin et al., 2019). We use 3 layers of transformer encoder, each with multi-head self-attention (=8), feed-forward layers with hidden dimension 4 and position embedding. The transformer computes attention across all input slots and masked slots , thus modeling dependencies and interactions between unmasked slots effectively. It results in assignments of temporally fused slots .
By leveraging temporal positional embeddings and masking, the temporal slot transformer achieves robust temporal alignment and enhances long-range temporal reasoning. In addition, it can also serve as a future slot prediction module.
3.4 Slot decoder
As our training objective is feature reconstruction of individual frames, we apply the decoder to all slots in at image level to obtain the reconstructed video features = . We experimented with two types of slot decoders.
MLP broadcast decoder
: A simple MLP broadcast decoder (Watters et al., 2019) has been used in VAE models and previous slot attention research (Seitzer et al., 2023). Each slot among the slots is broadcast to match the number of spatial patches resulting in tokens for each slot. A learnable positional encoding is added to each token. These tokens are then processed independently using a shared MLP to output reconstructed features and associated alpha masks indicating the slot’s attentive region. The final reconstruction is obtained by a weighted sum:
where denotes element-wise multiplication. The advantage of this simple design is its efficiency: as the MLP is shared across slots and positions, produced by the decoding is directly deployed as object segmentation masks.
SlotMixer decoder
: The recently introduced SlotMixer decoder (Sajjadi et al., 2022) for 3D object-centric learning has a constant overhead in the number of slots, as it only decodes once per output, requiring less computation. It employs an attention-based approach and operates in three key steps: slots allocation, mixing, and rendering (Sajjadi et al., 2022). The allocation step takes as input the slots and outputs an embedding vector using a cross-attention transformer. The mixing step is similar to a single-head attention step, using the embedding as queries and the slots as keys, to form an attention map , and the slot mix is obtained as
Finally the rendering step uses a MLP with position embedding shared with the allocation step to decode into , avoiding the need for a broadcast operation. The learned attention matrix is applied as channel segmentation masks with resolution .
Our experimental results demonstrate that, whether using an MLP broadcast decoder or a Slot-Mixer decoder, our method improves upon the vanilla RNN-based baseline and outperforms other state-of-the-art object-centric methods on the surgical video dataset.
3.5 Slot Contrastive Learning
To encourage diversity among slots and reduce redundancy, we adopt a contrastive learning loss based on cosine similarity, inspired by the SimCLR framework (Chen et al., 2020b). Unlike SimCLR, which computes similarity for positive pairs, our loss enhances dissimilarity between negative pairs (Radford et al., 2021). Let represent the one of slot vectors in for a given frame. For each slot vector , we compute its cosine similarity with all other slots within the same frame:
To exclude self-similarity, the cosine similarity matrix is adjusted by subtracting the identity matrix:
where is the Kronecker delta (i.e., if , and otherwise).
The total contrastive loss for each pair of slot vectors across frames is then computed as:
where is a temperature parameter. This loss function enhances slot diversity by enforcing the orthogonality between different slot vectors within each frame while preserving their magnitude.
3.6 Training loss
The main learning objective is to decode final slots to feature space . Reconstruction loss
guides the alignment between the original and reconstructed features.
The final loss function combines the reconstruction loss and the contrastive loss:
where is a scaling factor to balance the two terms.
The inclusion of the contrastive loss ensures that the slot representations remain diverse, while the reconstruction loss helps align the slots with the input features, guiding effective representation learning.
4 Experiments and Results
4.1 Dataset
We evaluate the performance of Slot-BERT on 4 surgical video dataset from 3 types of surgery (Table 1):
MICCAI 2022 SurgToolLoc Challenge Data: This dataset consists of 24,642 30 second video clips at 60 FPS collected from animal, phantom, and simulator-based surgeries (Zia et al., 2023). It includes a total of 13 instrument types, with up to four instruments in a clip. For simplicity, we refer to this dataset as MICCAI. For our use we downsample clips to 1 FPS.
Cholec80: The Cholec80 dataset (Twinanda et al., 2016) contains 80 cholecystectomy surgery videos downsampled to 1 FPS with presence labels for seven surgical tools in each frame. Following the preprocessing steps of a weakly supervised learning study (Liao et al., 2024), we derived a Transient Object Presence (TOP) video subset from Cholec80 resulting in 5,296 clips with 30 frames each. The test set is based on 100 clips from the CholecSeg8K dataset (Hong et al., 2020), which contains 8,000 frames with instrument and anatomy segmentation from 17 Cholec80 videos.
EndoVis 2017 Robotic Instrument Segmentation Challenge Data: This dataset contains 2,400 annotated frames from videos of abdominal surgery sampled at 1 FPS. We generated 480 video clips by grouping five consecutive frames per clip. The data includes seven instrument types with a maximum of four instruments present in a single clip.
Thoracic Robotic Surgery: This dataset includes data from 40 robot-assisted right upper lobectomies (RULs) for lung malignancy performed at the Toronto General Hospital between 2014 and 2023. A total of 264 annotated clips were selected for this study.
| MICCAI | Cholec | Endovis | Thoracic | |
| Procedure | abdominal | cholecystectomy | abdominal | thoracic |
| Number of clips | 24642 | 6300 | 480 | 264 |
| Frames per clip | 30 | 30 | 5 | 30 |
| Annotated clips | 100 | 100 | 480 | 264 |
| Annotated frames | 3000 | 3000 | 2400 | 550 |
| Clips in training set | 24542 | 5296 | — | — |
4.2 Metrics
We evaluate our approach based on the quality of the slot masks produced by the decoder using four primary metrics: video foreground ARI (FG-ARI) (Greff et al., 2019), video mean best overlap (mBO) (Pont-Tuset et al., 2016), mean best Hausdorff Distance (mBHD) and CorLoc(Bilen & Vedaldi, 2016) for evaluating localization accuracy. FG-ARI, adapted from a widely-used metric in object-centric research, measures the similarity between predicted object masks and ground truth masks, focusing on how effectively objects are segmented. mBO, on the other hand, evaluates the alignment between predicted and ground truth masks using the intersection-over-union (IoU) metric. For mBO, each ground truth mask is matched to the predicted mask with the highest IoU via Hungarian matching, and the average IoU is computed across all matched pairs. While FG-ARI primarily measures segmentation quality, mBO provides a more comprehensive assessment by including background pixels. Furthermore, the video version of mBO (mBO-V) also accounts for the temporal consistency of masks throughout the video. To facilitate comparisons at the frame level, we include the image-based version of mBO (mBO-F), which is computed on individual frames. This metric evaluates segmentation quality at the image level without considering temporal consistency. Additionally, to assess the boundary accuracy of the predicted masks, we calculate the mean best Hausdorff Distance (mBHD) for the best-overlapping masks produced by the object-centric learning model. CorLoc is also calculated based on overlapped object instance bounding box, given a threshould IoU0.5.
4.3 Baseline models
We compare our method with state-of-the-art object-centric algorithms, including SAVi (Kipf et al., 2022), STEVE (Singh et al., 2022), DINOSaur (Seitzer et al., 2023), Video-Saur (Zadaianchuk et al., 2023), and Slot-Diffusion (Wu et al., 2023b).
SAVi (Kipf et al., 2022) is a weakly supervised, recurrent video object-centric method that leverages the temporal dynamics of video data through optical flow. It conditions the initial slots on the central coordinates of objects. Since our training data lacks any form of supervision, we adopt a fully unsupervised version of SAVi that excludes initial state conditioning.111https://github.com/google-research/slot-attention-video/
STEVE (Singh et al., 2022) is an unsupervised model designed for object-centric learning in videos. Like SAVi, it employs an RNN-like slot initialization mechanism. However, STEVE replaces the broadcast decoder used in video slot attention with a specialized transformer-based slot decoder conditioned on slots. Its learning objective focuses on reconstructing individual video frames.222https://github.com/singhgautam/steve
DINOSaur (Seitzer et al., 2023) bridges the slot attention algorithm to real-world images by replacing the traditional image reconstruction objective with feature reconstruction. It can be implemented with feature encoders such as DINO, MAE, MoCo-v3 (Chen et al., 2021), or MSN (Assran et al., 2022), and the encoder can be either Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs). For our comparisons, we use the version based on the DINO ViT feature extractor.333https://github.com/amazon-science/object-centric-learning-framework
Video-Saur (Zadaianchuk et al., 2023), is a recurrent video object-centric method that replaces feature map reconstruction with next-frame feature cosine similarity as its reconstruction objective. It also employs a video-specific slot-mixer decoder for object-centric learning. Following the official implementation of Video-Saur, we report results using a combination of feature reconstruction and similarity reconstruction, with the DINO ViT as the feature extractor.444https://github.com/martius-lab/videosaur
Slot-Diffusion (Wu et al., 2023b) is an object-centric Latent Diffusion Model (LDM) designed for both image and video data. It introduces the replace of the decoding module for object-centric learning with slot-conditioned diffusion generation models. Based on the video version of Slot-Diffusion, similar to the learning objective of other diffusion models (Rombach et al., 2022), we trained it to reconstruct images of surgical videos following its encoding and decoding pipeline555https://github.com/Wuziyi616/SlotDiffusion.
4.4 Experiment setup
All raw images of video from different datasets were first cropped to remove zero pixels and then resized to 224224. For evaluating segmentation masks, they are upscaled to the same size for computing evaluation metrics. Our training and inference are implemented on a workstation with 503 GB ram based NVIDIA RTX 6000 Ada Generation GPUs. Adam optimizer with a learning rate of and a weight decay of was used for training. All models are trained under batch size of 4.
We first trained the models on the MICCAI and Cholec datasets. Training duration was 80 epochs for MICCAI and 100 epochs for Cholec. Performance of these models was then tested on all annotated frames from the MICCAI and Cholec testing sets. Next, we tested performance in an unsupervised transfer learning scenario. We used the MICCAI pre-trained weights as the original model, fine-tuned them on the Cholec training data for 10 epochs and measured performance on the Cholec testing set. To measure how well the learned object-centric representations generalize to novel, unseen databases without additional training we measured zero-shot segmentation performance by using the MICCAI trained models to segment testing sets in the EndoVis, Cholec, and Thoracic datasets.
All models were trained with the first 5 frames of each clip, even when the clips had longer sequences. For testing, each video was repeatedly processed by the trained model three times to calculate an average score and standard deviation, ensuring that any potential instabilities in the model’s outputs were accounted for (Wu et al., 2023b).
Finally, we tested performance on longer frame sequences of 7, 11, and 30 frames and evaluated a variation of our model on longer sequences by duplicating the TST module as a next-slot initializer.
4.5 Unsupervised segmentation performance
| Datasets | Method | mBO-V (%) | mBO-F (%) | mBHD (↓) | FG-ARI (%) | CorLoc (%) | |
| MICCAI | DINO-Saur(Seitzer et al., 2023) | 38.20 ± 0.90 | 42.90 ± 0.50 | 62.5 ± 1.3 | 48.40 ± 0.70 | 45.8 ± 2.2 | |
| SAVi(Kipf et al., 2022) | 29.40 ± 0.20 | 33.20 ± 0.10 | 81.7 ± 1.0 | 36.60 ± 0.20 | 40.0 ± 0.5 | ||
| STEVE(Singh et al., 2022) | 27.90 ± 0.20 | 31.50 ± 0.10 | 139.9 ± 0.6 | 34.30 ± 0.10 | 17.0 ± 0.4 | ||
| Slot-Diffusion(Wu et al., 2023b) | 37.50 ± 0.10 | 42.20 ± 0.10 | 70.5 ± 0.2 | 46.30 ± 0.00 | 42.0 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 46.30 ± 0.40 | 50.10 ± 0.40 | 53.9 ± 1.3 | 55.10 ± 0.50 | 60.1 ± 1.9 | ||
| Ours | 48.90 ± 0.20 | 52.80 ± 0.30 | 44.2 ± 0.6 | 58.20 ± 0.20 | 70.7 ± 0.8 | ||
| Cholec | DINO-Saur(Seitzer et al., 2023) | 25.70 ± 1.00 | 25.50 ± 0.50 | 75.9 ± 0.6 | 33.90 ± 0.70 | 29.1 ± 0.5 | |
| SAVi(Kipf et al., 2022) | 18.90 ± 0.20 | 18.10 ± 0.10 | 107.0 ± 0.2 | 23.70 ± 0.10 | 15.3 ± 0.1 | ||
| STEVE(Singh et al., 2022) | 19.50 ± 0.10 | 18.70 ± 0.00 | 108.9 ± 0.2 | 24.20 ± 0.00 | 19.8 ± 0.7 | ||
| Slot-Diffusion(Wu et al., 2023b) | 12.80 ± 0.00 | 15.70 ± 0.00 | 91.4 ± 0.3 | 19.40 ± 0.00 | 18.9 ± 0.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 26.10 ± 0.30 | 25.80 ± 0.20 | 74.9 ± 0.3 | 34.20 ± 0.20 | 34.6 ± 1.9 | ||
| Ours | 28.80 ± 0.30 | 27.80 ± 0.40 | 64.8 ± 0.9 | 37.00 ± 0.60 | 35.3 ± 2.0 |
This section presents the performance of models trained from scratch on MICCAI and Cholec data. Table 2 highlights the effectiveness of our approach across multiple evaluation metrics, demonstrating its robustness to various dataset characteristics.
All methods achieve their highest performance on the MICCAI dataset, reflecting the advantages of having a larger collection of data. Our method surpasses all competing methods in every metric: mBO-V (48.90%; +2.60% increase), mBO-F (52.80%; +2.70% increase), mBHD (44.2; -9.7 decrease), FG-ARI (58.20%; 3.10% increase) and CorLoc (70.7%; 10.6% increase). These improvements suggest that the model benefits from the dataset’s scale and diversity. Compared to the closest competitor, Video-Saur, our method significanly improves mBO-V, mBHD and CorLoc metrics, indicating better segment boundary and instance accuracy.
The performance on the Cholec dataset is comparatively lower for all models. Nevertheless, our method outperforms all baselines, with notable gains over Video-Saur (e.g., +2.7% in MBO-V, +2.8% in FG-ARI and -10.1 in mBHD). It is evident that the smaller size of the dataset and less diversity in the data impact performance compared to MICCAI, the latter being almost five times larger than Cholec. An additional challenge in Cholec dataset is the frequent instrument disappearance/reappearance (Liao et al., 2024) and the small size of some objects which pose a challenge because the slot number is fixed. These data characteristics also result in the phenomenon where mBO-F is smaller than mBO-V, unlike in MICCAI data, where most methods achieve a higher mBO-F instead.
4.6 Transfer learning performance
| Methods | mBO-V (%) | mBO-F (%) | FG-ARI (%) |
| DINO-Saur(Seitzer et al., 2023) | 26.3 ± 1.0 | 26.0 ± 0.9 | 34.7 ± 1.1 |
| SAVi(Kipf et al., 2022) | 18.2 ± 0.1 | 17.6 ± 0.0 | 23.0 ± 0.1 |
| STEVE(Singh et al., 2022) | 17.4 ± 0.0 | 16.7 ± 0.0 | 22.1 ± 0.1 |
| Slot-Diffusion(Wu et al., 2023b) | 23.4 ± 0.1 | 22.9 ± 0.1 | 29.8 ± 0.1 |
| Video-Saur(Zadaianchuk et al., 2023) | 28.7 ± 0.5 | 28.3 ± 0.5 | 37.3 ± 0.6 |
| Ours | 31.4 ± 0.6 | 30.1 ± 0.3 | 40.0 ± 0.7 |
We evaluate the transfer learning performance of Slot-BERT by taking the model trained on the MICCAI dataset and fine-tuning it on the Cholec dataset. We fine-tune the model for 10 epochs. Our experiments show that fine-tuning provides significant performance gains over training the model from scratch on the Cholec dataset indicating that object specific representations benefit from training on larger datasets and can be easily reused on smaller datasets.
A comparison of transfer learning performance with SOTA is summarized in Table 3. We observe that Slot-BERT outperforms all other methods in all three metrics, outperforming Video-Saur by 2.7% in mBO-V and FG-ARI. Fine-tuning results in an increase of 2.6% in mBO-V, 2.2% in mBO-F, and 3.0% in FG-ARI compared to training on Cholec data, confirming the plausibility of transfer learning to leverage large databases and enhance the performance of object-centric models on new datasets.
4.7 Zero-shot performance
Table 4 illustrates the performance of models trained on MICCAI and tested directly on unseen datasets without additional supervision. When comparing the zero-shot results to models trained fully from scratch on the respective testing domains (refer to Table 2), our method demonstrates performance that is highly competitive with domain-specific training, achieving almost identical results on the Cholec database and commendable results on the new EndoVis and Thoracic datasets.
Qualitative results of the zero-shot performance, illustrated in Figure 2, show that our method successfully segments unseen surgical instruments in new surgical scenes. In contrast, methods like STEVE and SAVi exhibit degraded adaptability, failing to capture objectiveness or accurately locate instruments in unseen videos.
As the Thoracic dataset is sparsely annotated (approximately two frames per clip), the mBO-V is less relevant than mBO-F. Looking at the results for the Thoracic dataset where frame lengths differ significantly compared to Endovis and Cholec, our model still outperforms others, achieving an 2.2% better mBO-F and 6.4 decrease in mBHD compared to Video-Saur. These results validate the robustness of our model in diverse settings.
Figure 2 also reveals a drawback of our approach for the accuracy of segmentation masks. As the frames are divided into patches, our method is not accurate at detecting precise object boundaries, in particular for instruments, as the immediate background of instruments is highly correlated across scenes. The method tends to however correctly locate the whole object, leading to a low false negative rate.
| Dataset | Method | mBO-V (%) | mBO-F (%) | mBHD (↓) | FG-ARI (%) |
| Endovis | DINO-Saur(Seitzer et al., 2023) | 34.1 ± 0.3 | 38.7 ± 0.4 | 64.4 ± 1.1 | 45.4 ± 0.5 |
| SAVi(Kipf et al., 2022) | 28.2 ± 0.2 | 32.4 ± 0.1 | 81.6 ± 0.2 | 36.9 ± 0.0 | |
| STEVE(Singh et al., 2022) | 26.0 ± 0.1 | 29.8 ± 0.1 | 143.7 ± 0.2 | 33.9 ± 0.1 | |
| Slot-Diffusion(Wu et al., 2023b) | 33.3 ± 0.1 | 35.9 ± 0.1 | 95.6 ± 0.1 | 40.6 ± 0.1 | |
| Video-Saur(Zadaianchuk et al., 2023) | 42.7 ± 0.2 | 46.7 ± 0.2 | 56.2 ± 0.2 | 53.5 ± 0.1 | |
| Ours | 43.5 ± 0.3 | 47.6 ± 0.3 | 50.7 ± 0.4 | 54.4 ± 0.3 | |
| Thoracic † | DINO-Saur(Seitzer et al., 2023) | 27.7 ± 0.6 | 34.7 ± 0.5 | 90.7 ± 0.7 | 28.2 ± 0.5 |
| SAVi(Kipf et al., 2022) | 24.9 ± 0.2 | 28.5 ± 0.0 | 105.3 ± 1.0 | 21.7 ± 0.2 | |
| STEVE(Singh et al., 2022) | 24.0 ± 0.0 | 30.2 ± 0.1 | 127.9 ± 0.6 | 22.6 ± 0.0 | |
| Slot-Diffusion(Wu et al., 2023b) | 29.3 ± 0.1 | 36.9 ± 0.1 | 103.1 ± 0.2 | 28.1 ± 0.1 | |
| Video-Saur(Zadaianchuk et al., 2023) | 36.7 ± 0.1 | 48.9 ± 0.1 | 72.8 ± 0.4 | 39.2 ± 0.1 | |
| Ours | 37.7 ± 0.2 | 51.1 ± 0.3 | 66.4 ± 0.3 | 40.5 ± 0.1 | |
| Cholec | DINO-Saur(Seitzer et al., 2023) | 21.5 ± 0.8 | 20.7 ± 0.4 | 82.8 ± 1.3 | 28.1 ± 0.9 |
| SAVi(Kipf et al., 2022) | 17.8 ± 0.1 | 17.3 ± 0.1 | 106.1 ± 0.7 | 22.6 ± 0.1 | |
| STEVE(Singh et al., 2022) | 17.3 ± 0.1 | 16.6 ± 0.1 | 139.2 ± 0.9 | 21.9 ± 0.1 | |
| Slot-Diffusion(Wu et al., 2023b) | 17.8 ± 0.1 | 17.0 ± 0.0 | 106.1 ± 0.3 | 20.9 ± 0.1 | |
| Video-Saur(Zadaianchuk et al., 2023) | 27.0 ± 0.5 | 26.1 ± 0.5 | 72.5 ± 1.5 | 35.5 ± 0.6 | |
| Ours | 29.2 ± 0.1 | 27.9 ± 0.1 | 64.0 ± 0.6 | 37.7 ± 0.0 |
†When zero-shot on thoracic the video length is increased to 30 frames in comparison to Endovis and Cholec (5 frames).
4.8 Adaptation to longer sequences
| 7 Frames | 11 Frames | ||||||||
| Method | mBO-V (%) | mBO-F (%) | mBHD (↓) | FG-ARI (%) | mBO-V (%) | mBO-F (%) | mBHD (↓) | FG-ARI (%) | |
| DINO-Saur(Seitzer et al., 2023) | 33.50 ± 0.90 | 39.80 ± 0.20 | 64.0 ± 1.9 | 44.90 ± 0.30 | 32.00 ± 0.30 | 40.20 ± 0.40 | 62.8 ± 1.1 | 44.30 ± 0.40 | |
| SAVi(Kipf et al., 2022) | 28.80 ± 0.30 | 33.50 ± 0.20 | 80.9 ± 0.9 | 36.90 ± 0.20 | 27.40 ± 0.20 | 33.70 ± 0.30 | 81.7 ± 0.5 | 36.00 ± 0.30 | |
| STEVE(Singh et al., 2022) | 27.10 ± 0.10 | 31.40 ± 0.10 | 141.0 ± 0.7 | 34.30 ± 0.10 | 25.70 ± 0.00 | 31.40 ± 0.00 | 140.3 ± 0.7 | 33.40 ± 0.00 | |
| Slot-Diffusion(Wu et al., 2023b) | 35.60 ± 0.20 | 38.30 ± 0.10 | 79.6 ± 0.1 | 41.20 ± 0.10 | 34.00 ± 0.10 | 38.20 ± 0.10 | 80.2 ± 0.2 | 40.00 ± 0.10 | |
| Video-Saur(Zadaianchuk et al., 2023) | 46.70 ± 0.30 | 51.10 ± 0.20 | 52.5 ± 0.9 | 56.10 ± 0.30 | 44.40 ± 0.30 | 51.60 ± 0.20 | 51.5 ± 0.4 | 55.20 ± 0.20 | |
| Ours | 48.00 ± 0.50 | 52.30 ± 0.30 | 44.1 ± 0.7 | 57.50 ± 0.50 | 46.20 ± 0.50 | 53.10 ± 0.30 | 43.8 ± 0.9 | 56.80 ± 0.40 | |
| Ours + Future slot prediction | 48.50 ± 0.80 | 52.80 ± 0.70 | 44.14 ± 1.1 | 58.00 ± 0.90 | 46.90 ± 0.00 | 53.60 ± 0.10 | 43.10 ± 0.16 | 57.30 ± 0.30 | |
To evaluate the robustness and generalization of our method to longer video sequences, we tested our model on sequences of 7 and 11 frames after training on 5 frames. A sliding window approach was employed to handle longer inputs, ensuring temporal alignment during prediction. Additionally, we introduced a future slot prediction mechanism that replaced the simple RNN-based slot initialization which is only possible when testing on video length that is equal to length in training. In this new design, the temporal slot transformer is fed all previous slots, leaving the slot at the last position empty. The temporal slot transformer then predicts this missing slot, which serves as the initialization for subsequent frames. An illustration of this next slot initialization can be found in supplementary material section A.
The results in Table 5 demonstrate that our method achieves better performance compared to state-of-the-art approaches across all metrics, particularly in terms of mBO-V, mBO-F, MBHD and FG-ARI. For instance, in 7-frame sequences, our method achieved an mBO-V of 48.0% and an FG-ARI of 57.5%, surpassing the next best method, Video-Saur, by 1.3% and 1.4%, respectively. When extended to 11-frame sequences, our method maintained its robust performance, achieving an mBO-V of 46.2% and an FG-ARI of 56.8%. These results highlight the minimal degradation in accuracy as the sequence length increases, demonstrating the adaptability of our approach.
With the future slot prediction mechanism, we observed further performance gains, particularly for longer sequences with 11 frames. The mBO-V increased to 46.9%, and the FG-ARI improved to 57.3%, highlighting the effectiveness of this advanced initialization strategy in maintaining temporal coherence.
Figure 3 visualizes the impact of sequence length on accuracy. Across methods, mBO-F showed minimal variation with longer sequences, demonstrating that frame-level accuracy is generally stable across sequence lengths. However, the challenge of maintaining temporal consistency is evident in video-level metrics (mBO-V). Most methods, including SAVi, STEVE, and Video-Saur, exhibited a notable decline in mBO-V with longer sequences. In contrast, our method maintains the highest mBO-V among all evaluated methods. This robustness underscores our approach’s ability to adapt to varying temporal contexts effectively.
| Same domain (MICCAI) | Zero shot transfer | ||||||||||
| Short (5 frames) | Long (11 frames) | Endovis | Thoracic | ||||||||
| Method | mBO-V (%) | FG-ARI (%) | mBO-V (%) | FG-ARI (%) | mBO-V (%) | FG-ARI (%) | mBO-V (%) | FG-ARI (%) | |||
| w/o contrast w/o TST | 45.0 ± 0.5 | 53.5 ± 0.5 | 42.0 ± 0.5 | 53.2 ± 0.6 | 39.5 ± 0.2 | 50.1 ± 0.2 | 34.5 ± 0.1 | 36.6 ± 0.1 | |||
| contrast only | 46.4 ± 0.6 | 55.3 ± 0.6 | 44.1 ± 0.1 | 55.0 ± 0.2 | 42.1 ± 0.4 | 52.6 ± 0.3 | 34.2 ± 0.1 | 37.5 ± 0.2 | |||
| w/o contrast | 46.3 ± 0.4 | 55.2 ± 0.4 | 43.7 ± 0.3 | 54.7 ± 0.1 | 40.8 ± 0.3 | 51.5 ± 0.2 | 34.1 ± 0.2 | 36.7 ± 0.2 | |||
| maskout feature | 45.8 ± 0.5 | 55.0 ± 0.5 | 42.2 ± 0.6 | 53.6 ± 0.7 | 41.0 ± 0.5 | 51.7 ± 0.3 | 32.9 ± 0.2 | 36.6 ± 0.1 | |||
| w/o slot masks | 48.6 ± 0.3 | 57.6 ± 0.5 | 45.9 ± 0.9 | 57.1 ± 0.4 | 43.0 ± 0.3 | 53.9 ± 0.2 | 35.0 ± 0.1 | 38.7 ± 0.1 | |||
| our full | 48.9 ± 0.2 | 58.2 ± 0.3 | 46.9 ± 0.0 | 57.3 ± 0.3 | 44.0 ± 0.2 | 55.2 ± 0.3 | 37.7 ± 0.2 | 40.5 ± 0.1 | |||
4.9 Ablation study
To investigate the contributions of individual components in our model to its final performance, we systematically removed or modified different components and observed their impact on segmentation accuracy, as summarized in Table 6. These experiments were conducted in both same-domain and zero-shot transfer scenarios, with evaluations on the MICCAI dataset (short and long sequences), EndoVis dataset, and Thoracic dataset. Below, we discuss the implications of the results for each configuration.
Baseline: In this configuration, both the contrastive loss and TST module were removed, and a simple recurrent refinement network (RRN) was employed along with a standard four-layer MLP decoder (denoted as w/o contrast w/o TST). This baseline yielded the lowest performance across all datasets and settings. For instance, on the MICCAI test set (long sequences), the mBO-V score dropped to , while on EndoVis, it fell to . These results highlight the critical contributions of both the contrastive loss and the TST module in enhancing segmentation quality.
Slot Contrast only: Removing the slot-BERT module while retaining the contrastive loss (denoted as contrast only) resulted in improvements over the vanilla baseline. For instance, on the MICCAI dataset (short sequences), the mBO-V score increased to , and the FG-ARI improved to . However, the performance remained below that of the full model, underscoring the role of slot-BERT in refining slot representations and contextualizing features effectively.
Effect of slot contrastive loss: In this configuration, the contrastive loss was omitted while retaining the slot-BERT module (denoted as w/o contrast). The results were similar to those of the previous configuration, with slight decreases in performance on most datasets. For example, on the EndoVis dataset, the mBO-V score was , which is lower than the configuration without TST but with contrastive loss. These findings indicate that the contrastive loss is pivotal for improving slot distinctiveness, particularly in datasets with diverse instrument appearances.
Effect of slot masks: We also examined the impact of slot-specific masking by directly inputting video slots into the TST module without applying slot masks and replacing the masked transformer with a standard transformer (referred to as w/o slot masks). This led to a slight performance degradation. For instance, on the MICCAI dataset (long sequences), the mBO-V score was compared to for the full model. This suggests that slot masks do contribute to robust representation learning by focusing on relevant feature regions.
In addition, we replaced slot-specific masking with feature-level masking. A masked autoencoding strategy was applied, where random feature patches were masked during training (denoted as maskout feature). This configuration showed further degradation in performance compared to the full model. For example, on the Thoracic dataset, the mBO-V score was , significantly lower than the achieved by the full model. This indicates that random feature masking is less effective than the structured masking strategy employed by the slot-BERT module.
The results in Table 6 demonstrate the critical role of each component in our model. The full model consistently outperformed all other configurations across datasets and metrics. Notably, it achieved the highest mBO-V scores in all settings, including on MICCAI (short sequences), on EndoVis, and on Thoracic. These results validate the synergistic contributions of the TST module, contrastive loss, and slot-specific masking in achieving superior segmentation performance. Specifically, the contrastive loss enhances inter-slot separation, the TST module improves feature contextualization and representation, and slot-specific masking ensures better focus on relevant regions. The full model’s superior performance across both same-domain and zero-shot scenarios underscores the importance of integrating these components for robust segmentation in diverse domains.
4.10 Experiment with slot-Mixer Decoder
| Setup | MICCAI | |||
| mBO-V | mBO-F | mBHD() | CorLoc | |
| Mixer only | 47.1±0.4 | 51.2±0.3 | 50.889±0.326 | 61.6±0.5 |
| Contrast+Mixer | 48.9±0.6 | 53.1±0.5 | 48.121±1.197 | 66.2±0.7 |
| Mixer-full | 49.0±0.4 | 53.2±0.2 | 46.994±1.008 | 67.4±0.9 |
| Setup | Zero shot to Endovis | |||
| mBO-V | mBO-F | mBHD () | CorLoc | |
| Mixer only | 42.7±0.3 | 47.1±0.2 | 54.321±0.6 | 58.1±0.6 |
| Contrast+Mixer | 43.3±0.1 | 47.5±0.2 | 53.021±0.279 | 60.9±1.1 |
| Mixer-full | 43.2±0.2 | 47.5±0.2 | 51.945±0.5 | 61.9±0.3 |
To evaluate our method with the alternative Slot-Mixer decoder (Zadaianchuk et al., 2023), we conducted experiments using the Slot-BERT framework and compared it against two additional configurations. The results are reported for the MICCAI benchmark, where the model is trained and tested on MICCAI, and the zero-shot Endovis benchmark, where the model trained on MICCAI is tested directly on Endovis without fine-tuning. The first configuration, Mixer Only, combines a vanilla RNN Slot Attention with the Slot-Mixer for feature reconstruction. The second, Contrast+Mixer, incorporates slot contrastive loss into the training process while still relying on the Slot-Mixer for decoding. Finally, Mixer-full represents the full proposed method.
As shown in table 7, in the MICCAI benchmark, the proposed method achieves an mBHD error of 46.9, improving upon the results of baseline of Mixer-only 50.8. For Endovis, in zero-shot transfer, our method achieves mBHD of 51.9. This demonstrates the comparable performance of the Slot-Mixer decoder to the MLP decoder:44.2 mBHD on MICCAI and 50.7 mBHD zero shot to Endovis (table 2 and table 4).
These findings confirm that our method is able to surpass the vanilla RNN-based baseline using either MLP broadcast or Slot-Mixer decoder, achieving high localization and segmentation performance. For more qualitative results refer to the section F.1 of supplementary materials.
4.11 Affect of the amount of training data on performance
We evaluate the sensitivity of segmentation performance to the amount of training data. Figure 4 compares Slot-BERT, our proposed method, with Video-Saur on the MICCAI challenge dataset, using mBO-V as video-level segmentation accuracy evaluation metrics. Training data proportions range from the full dataset (1.0) to 1% (0.01).
Slot-BERT demonstrates remarkable robustness in dealing with a decrease in training data. With the full dataset, it achieves 48.9% accuracy compared to 46.3% for Video-Saur. With only 30% of the data, Slot-BERT maintains 47.5%, nearly matching its full training data performance, while Video-Saur drops to 44.4%. At lower data ratios, the gap widens significantly: Slot-BERT achieves 30.4% at 0.05 compared to Video-Saur’s 7.3%, and at 0.01, Slot-BERT retains 19.6% while Video-Saur falls to 6.0%. In summary, Slot-BERT consistently outperforms Video-Saur across all data proportions and maintains higher accuracy even with minimal data. Its graceful degradation in performance highlights its robustness and makes it particularly effective for segmentation tasks in data-scarce scenarios.
4.12 Experiment with different slot numbers
In order to investigate the effect of changing slot number on the performance of Slot-BERT, experiments were conducted using MLP and Mixer decoders. Figure 5 presents the CorLoc score obtained on training domain and zeroshot data, to evaluate impact of slot allocation on the localization accuracy. The results show that generally provides the best performance for both methods, with MLP achieving a peak score of 70.7 and Mixer reaching 67.4 on training domain data. As increases beyond 7, performance begins to decline for both decoders, suggesting that excessively large values might over-delineate the image as the number of objects in a video is limited. MLP consistently outperforms Mixer on training domain data, indicating its superior ability to leverage increased for improved predictions.
On zero-shot data, the trend remains consistent, with both MLP and Mixer achieving their best scores at , though their overall performance is lower compared to the training domain. MLP slightly outperforms Mixer across most values, reaching a maximum zero-shot score of 62.8, compared to Mixer’s 61.9 at . These results highlight that while is optimal across both datasets, there is a clear performance gap between training domain and zeroshot data, reflecting challenges in generalization. The analysis demonstrates that adjusting significantly impacts the effectiveness of both decoding methods, emphasizing the importance of finding an optimal value to balance performance and generalization. Evaluation of different slot number under additional metrics are concluded in Supplementary Material section D.
5 Conclusions
Slot-BERT demonstrated robust spatial segmentation and temporal reasoning capabilities, achieving state-of-the-art results across diverse datasets. The proposed bidirectional slots transformer enabled reasoning over video sequences without the scalability issues typical of parallel methods, because the attention operates on lightweight slot embeddings. This capability is particularly relevant for domains requiring long-term temporal coherence, such as surgical video analysis.
By enforcing orthogonality between slots, our contrastive loss improved the independence of latent representations, reducing leakage across slot boundaries. This enhancement translated into more precise segmentation maps and improved object discovery, even in zero-shot settings.
Unlike existing models that either simply update slots recursively, lacking long-range reasoning, or process entire video features for grouping, which is computationally costly, Slot-BERT scales efficiently for videos. It achieves competitive performance without relying on computationally intensive modalities like optical flow or depth maps. This characteristic makes Slot-BERT a practical choice for real-world applications requiring affordable computational resources.
5.1 Limitations
The effectiveness of slot-based representations partially depends on the quality of initial slot assignments. Suboptimal initialization can impact downstream temporal reasoning and segmentation quality. The number of slots also plays a role in how well slots encompass particular object instances. The optimal number of slots, as well as dynamic slot allocation are part of current research (Fan et al., 2024).
Although effective in handling moderate temporal dynamics, scenarios involving rapid, non-linear object motion or occlusions may still present challenges for Slot-BERT.
While our model outperformed baselines in zero-shot generalization tasks, there is room for improvement in handling novel object classes or extreme domain shifts.
While our model discovers the overall location and shape of objects, it fails to predict the exact pixel-level boundaries, lowering the semantic segmentation accuracy. This is partially due to patch-based processing and could be mitigated by increasing the resolution of the patch grid. The model trained on lower-resolution path features can serve as guidance for higher-resolution modalities, such as optical flow and image saliency maps, enabling accurate high-resolution segmentation in an unsupervised manner. With Slot-BERT the stability of integrating these modalities will enhanced.
5.2 Future Directions
Incorporating additional priors or augmentations for slot initialization, such as unsupervised identifiable slot attention mechanisms (Kori et al., 2024), may further enhance segmentation and temporal reasoning capabilities. Slot attention is a fully un-supervised representation learning algorithm; however, it can also be implemented in weakly supervised learning or in a human-interactive manner to identify the class of each slot so that pseudo-segmentation masks can be generated. Specifically, with weak video class labels, set prediction frameworks can be adopted to the latent space of slots (Sun et al., 2021). Alternatively, with pre-trained slot models, annotators only need to identify which slot needs to be tracked or which slot represents the object of interest. In such interactive implementations, Slot-BERT can reduce the annotation burden in labeling tasks or directly serve as an attention aid in tasks like surgical tracking and planning.
While Slot-BERT eschews computationally intensive modalities like optical flow or feature similarity, integrating lightweight versions of these signals could provide complementary information in challenging scenarios. Such modalities could be an option to increase the resolution of slot attention masks, which are currently limited by the resolution of patches. For instance, the similarity of images or lower-level features often has much higher resolution compared to high-level features.
Extending Slot-BERT to other biomedical domains, such as surgical videos involving cell segmentation in microscopic volumes or organ segmentation in MRI scanning sequences, would help evaluate its adaptability to diverse tasks.
Slot-BERT represents a significant advancement in object-centric video representation learning, addressing critical challenges in long video sequences of surgical domains. By integrating slot-based reasoning with bidirectional transformers, the model balances scalability and segmentation accuracy, paving the way for future innovations in self-supervised learning. Its demonstrated potential for zero-shot generalization highlights the importance of modular, explainable architectures for tackling real-world, domain-specific challenges.
Acknowledgments
This work was supported by the University of Pennsylvania Thomas B. McCabe and Jeannette E. Laws McCabe Fellow Award and the Linda Pechenik Montague Investigator Award.
Supplementary Material
A Illustration of using TST module for next slot initialization
In Section 4.8 of the main text, we present the results of adapting pre-trained slot-BERT for longer sequence prediction. In addition to using a traditional approach for slot initialization, where the previously predicted slot is fed into the next slot encoder as an initializer, we also demonstrate an alternative option of using the same TST module as a slot initializer, as shown in Figure S1. In this new design, we take the previous slot buffer and append an empty slot (zero vectors) to the latest position. The only modification to the TST module is to switch it from random masking to fixing the mask at the last location, enabling the TST module to predict the missing empty slot by reasoning over the historical slots. Finally, the next slot encoder takes the predicted slot from the TST module, , as the initialization to update slot . In our experiment, instead of creating a new TST module for the initializer or freezing the pre-trained TST for postprocessing, we allow the initializer and slot decoder to share weights from the same TST module and enable end-to-end training with the new slot initializer.
B Experiment with different slot contrastive temperatures
We conducted experiments to evaluate the impact of the slot contrastive temperature () in the calculation of contrastive loss on segmentation performance in both same-domain and zero-shot transfer scenarios. The model was trained on the MICCAI dataset and tested on both the MICCAI test set (same-domain performance) and the unseen EndoVis dataset (zero-shot transfer performance).
The results, presented in Table S1, demonstrate the nuanced role of contrastive temperature in influencing segmentation accuracy. In the same-domain context, a temperature setting of achieved the best mBO-V score of 49.6%, along with a frame-level accuracy (mBO-F) of 52.8%. Meanwhile, resulted in the highest mBO-F score on the MICCAI data (53.0%). This suggests that a smaller temperature effectively balances inter-slot distinctiveness while maintaining smooth transitions across highly similar instrument categories, which is a common characteristic of the MICCAI dataset, where some instruments share significant visual similarities.
In the zero-shot transfer setting tested on the EndoVis dataset, the performance trends shifted. Here, achieved the highest mBO-V score of 44.0% and an mBO-F of 48.4%. This indicates that a slightly higher temperature helps the model better separate and distinguish slots in a dataset with more varied instrument appearances, improving generalization. Interestingly, while yielded comparable performance in the zero-shot scenario (mBO-V: 43.6%, mBO-F: 47.5%), it decreased the segmentation accuracy compared to , suggesting that overly increasing the temperature can negatively impact performance.
These findings emphasize the importance of carefully selecting the contrastive temperature based on the target domain and task requirements. The contrastive temperature provides flexibility, enabling the model to adjust its behavior for optimal performance in both same-domain and zero-shot settings.
| Same domain | Zeroshot transfer | ||||
| mBO-V (%) | mBO-F (%) | mBO-V (%) | mBO-F (%) | ||
| 1.2 | 48.5 ± 0.5 | 52.2 ± 0.3 | 43.6 ± 0.4 | 47.5 ± 0.3 | |
| 1.1 | 48.9 ± 0.2 | 52.8 ± 0.2 | 44.0 ± 0.2 | 48.4 ± 0.1 | |
| 1.0 | 48.2 ± 0.2 | 51.9 ± 0.3 | 43.5 ± 0.3 | 47.6 ± 0.3 | |
| 0.8 | 49.6 ± 0.1 | 52.8 ± 0.3 | 43.4 ± 0.3 | 47.3 ± 0.2 | |
| 0.5 | 48.4 ± 0.3 | 53.0 ± 0.2 | 43.0 ± 0.2 | 47.7 ± 0.1 | |
| 0.1 | 48.9 ± 0.5 | 52.6 ± 0.7 | 44.0 ± 0.2 | 47.5 ± 0.2 | |
C Visualization of latent slot embedding projection
To investigate the distinctiveness and relationships among slot vectors, we first visualize the latent slot embeddings learned by Slot-BERT using t-SNE (Van der Maaten & Hinton, 2008). As shown in Figure S2, slot vectors corresponding to different instruments within a video tend to cluster closer to each other, while remaining distinct from the clusters of tissue-related slot vectors. This behavior highlights the disentangled and interpretable nature of the slot representations learned by our approach.
To further validate this observation, we applied Principal Component Analysis (PCA), a linear projection method that preserves the relative distances between data points in the latent space. The PCA mapping also reveals clear separation between regions corresponding to instrument slots and tissue slots, while maintaining the internal distinctiveness of each category. These results indicate that the slot embeddings learned by Slot-BERT are both robust and explainable, effectively capturing the compositional structure of objects in video data.
D Results of different slot numbers with additional metrics
| Training Domain | Zero-Shot Domain | |||||||||||
| Decoder | mBO-V (%) | mBO-F (%) | mBHD () | FG-ARI (%) | CorLoc (%) | mBO-V (%) | mBO-F (%) | mBHD () | FG-ARI (%) | CorLoc (%) | ||
| MLP | 3 | 37.8 ± 0.3 | 40.5 ± 0.2 | 101.46 ± 0.93 | 42.3 ± 0.3 | 52.1 ± 0.4 | 33.4 ± 0.2 | 35.7 ± 0.1 | 100.59 ± 0.33 | 38.8 ± 0.1 | 49.2 ± 0.5 | |
| 5 | 45.6 ± 0.4 | 48.9 ± 0.5 | 53.25 ± 2.21 | 53.1 ± 0.6 | 66.2 ± 0.8 | 40.8 ± 0.1 | 44.2 ± 0.1 | 61.10 ± 0.35 | 49.7 ± 0.1 | 61.9 ± 0.5 | ||
| 7 | 48.9 ± 0.2 | 52.8 ± 0.2 | 43.40 ± 0.53 | 58.2 ± 0.3 | 70.7 ± 0.8 | 44.0 ± 0.2 | 48.4 ± 0.1 | 49.26 ± 0.93 | 55.2 ± 0.3 | 62.8 ± 0.2 | ||
| 9 | 49.1 ± 0.5 | 54.3 ± 0.1 | 41.81 ± 0.94 | 60.1 ± 0.2 | 69.0 ± 0.5 | 44.4 ± 0.1 | 49.4 ± 0.2 | 49.43 ± 0.18 | 56.9 ± 0.3 | 59.4 ± 0.5 | ||
| 11 | 49.2 ± 0.3 | 53.6 ± 0.1 | 44.47 ± 0.19 | 60.0 ± 0.2 | 63.7 ± 0.8 | 44.6 ± 0.4 | 48.9 ± 0.1 | 50.82 ± 0.43 | 57.1 ± 0.2 | 51.2 ± 0.5 | ||
| Mixer | 3 | 33.9 ± 0.4 | 36.3 ± 0.3 | 114.97 ± 1.23 | 36.2 ± 0.4 | 43.9 ± 0.8 | 30.0 ± 0.1 | 32.5 ± 0.1 | 110.21 ± 0.36 | 34.1 ± 0.1 | 39.6 ± 0.1 | |
| 5 | 45.2 ± 0.0 | 48.9 ± 0.1 | 59.85 ± 0.57 | 52.6 ± 0.2 | 62.3 ± 1.0 | 40.2 ± 0.2 | 43.6 ± 0.2 | 63.13 ± 0.60 | 49.0 ± 0.1 | 59.7 ± 0.2 | ||
| 7 | 49.0 ± 0.4 | 53.2 ± 0.2 | 46.99 ± 1.01 | 58.2 ± 0.2 | 67.4 ± 0.9 | 43.2 ± 0.2 | 47.5 ± 0.2 | 51.95 ± 0.50 | 54.3 ± 0.1 | 61.9 ± 0.3 | ||
| 9 | 49.3 ± 0.9 | 53.7 ± 0.7 | 46.41 ± 0.98 | 59.7 ± 0.7 | 61.9 ± 1.0 | 44.0 ± 0.3 | 49.1 ± 0.1 | 51.03 ± 0.18 | 56.6 ± 0.1 | 57.5 ± 0.5 | ||
| 11 | 47.7 ± 0.6 | 52.5 ± 0.2 | 49.73 ± 0.62 | 59.1 ± 0.1 | 55.6 ± 0.6 | 43.7 ± 0.3 | 48.8 ± 0.2 | 53.14 ± 0.65 | 57.1 ± 0.3 | 48.7 ± 0.5 | ||
| Transfer learning | From scratch | |||||||||
| Dataset | Method | mBO-V(%) | mBO-F(%) | mBHD (↓) | FG-ARI(%) | mBO-V(%) | mBO-F(%) | mBHD (↓) | FG-ARI(%) | |
| Endovis | DINO-Saur(Seitzer et al., 2023) | 38.8 | 44.1 | 52.4 | 51.3 | 25.6 | 26.1 | 122.5 | 29.4 | |
| SAVi(Kipf et al., 2022) | 30.3 | 34.1 | 77.8 | 39.3 | 34.4 | 38.5 | 110.9 | 42.9 | ||
| STEVE(Singh et al., 2022) | 26.4 | 30.3 | 147.2 | 34.5 | 31.5 | 34.1 | 97.9 | 38.6 | ||
| Slot-Diffusion(Wu et al., 2023b) | 36.2 | 38.5 | 66.7 | 43.3 | 42.3 | 46.0 | 56.9 | 53.2 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 46.9 | 51.2 | 50.8 | 57.8 | 25.7 | 25.8 | 135.0 | 39.4 | ||
| Ours | 48.8 | 52.3 | 41.7 | 59.2 | 45.1 | 46.0 | 70.1 | 51.6 | ||
| Thoracic | DINO-Saur(Seitzer et al., 2023) | 31.1 | 39.3 | 86.0 | 32.5 | 27.1 | 30.7 | 104.8 | 23.9 | |
| SAVi(Kipf et al., 2022) | 25.6 | 28.7 | 106.5 | 22.5 | 26.5 | 30.7 | 119.8 | 23.1 | ||
| STEVE(Singh et al., 2022) | 23.8 | 30.0 | 131.9 | 22.3 | 28.9 | 33.0 | 135.4 | 26.5 | ||
| Slot-Diffusion(Wu et al., 2023b) | 29.9 | 37.8 | 104.3 | 28.8 | 31.1 | 39.9 | 84.6 | 31.4 | ||
| Video-Saur(Zadaianchuk et al., 2023) | 38.9 | 52.1 | 65.7 | 41.9 | 21.9 | 15.7 | 139.5 | 11.9 | ||
| Ours | 39.9 | 52.1 | 65.2 | 41.7 | 34.0 | 40.5 | 92.4 | 33.8 | ||
In section 4.12 we present the results of our method using different slot numbers. And demonstrated that with slot number 7 it shows optimal localization accuracy when either MLP or slot-Mixer decoder is adopted for Slot-BERT. Here detailed quantitative results on additional metrics is revealed in table S2. As shown in Table S2, specifically, with 7 slots and an MLP decoder, the CorLoc metric reaches its peak at 70.7 ± 0.8 in the training domain, and mBHD achieves its value of 43.40 ± 0.53, reflecting accurate boundary localization. Similarly, with the Slot-Mixer decoder and 7 slots, the CorLoc metric achieves its highest value of 67.4 ± 0.9, demonstrating robust spatial localization.
When extending the analysis to evaluate segmentation overlap on the training and zero-shot domains, using 9 slots emerges as a strong performer. With the MLP decoder, 9 slots achieve the highest mBO-F at 54.3 ± 0.1 in the training domain and 49.4 ± 0.2 in the zero-shot domain, Moreover, the Slot-Mixer decoder with 9 slots delivers the best mBO-V scores of 49.3 ± 0.9 for the training and zero-shot domains, respectively, signifying enhanced temporal coherence in video object segmentation.
These results underscore the nuanced trade-offs across different slot numbers and decoder configurations. While slot number 7 provides strong spatial localization, slot number 9 excels in overlap accuracy and temporal consistency, making it a versatile choice depending on the task emphasis.
E Additional results on transfer learning on Endovis and Thoracic data
We fine-tuned the MICCAI pre-trained model on the EndoVis and Thoracic datasets for 80 epochs as an extention of the transfer learning experiment. Additionally, we trained all models from scratch on the EndoVis and Thoracic datasets for 2000 epochs, to compare their performance with transfer learning. To effectively utilize the small EndoVis and Thoracic datasets for fine-tuning, following previous unsupervised segmentation approaches (Zhang et al., 2015; Wang et al., 2017), we allowed the model access to the full EndoVis and Thoracic videos. Our goal was to evaluate whether the models could generate accurate pseudo-masks under the condition of no ground truth supervision signals.
As presented in Table S3, after fine-tuning on these two domain-specific datasets (EndoVis and Thoracic) using only self-supervised objectives such as reconstruction and contrastive learning, without relying on any labels, our method further improves performance. Remarkably, the fine-tuned results surpass those of models trained from scratch. For example, in EndoVis, our fine-tuned model achieves an FG-ARI of 59.2%, which is higher than the scratch-trained model’s FG-ARI of 51.6%. Similarly, on the Thoracic dataset, our method achieves notable gains in mBO-V and FG-ARI after fine-tuning, demonstrating the effectiveness of leveraging knowledge obtained from unsupervised pretraining on MICCAI data. It is worth noting that for some methods (e.g., Slot Diffusion and STEVE), transfer learning did not outperform training from scratch. This could possibly be because these methods tend to overfit on small datasets rather than generalize and effectively leverage knowledge across domains.
F Additional qualitative results
F.1 Comparison on using MLP and slot-Mixer decoder
In Section 4.10, we present the results of using the alternative Slot-Mixer decoder instead of the MLP decoder (as explained in Section 3.5) to reconstruct target features, exploring various setup variations. Detailed qualitative comparisons between different decoders are provided here. Figure S3 illustrates segmentation masks produced by Slot-BERT with both MLP and Slot-Mixer decoders, tested on sequences of 30 frames at 1 FPS from the MICCAI dataset. The model was trained using 7 slots.
Both MLP and Slot-Mixer decoders demonstrate unique strengths. With the MLP decoder, the segmentation masks tend to cover objects more comprehensively and accurately locate them. On the other hand, the Slot-Mixer decoder exhibits superior temporal consistency. For instance, in the first and third rows of Figure S3, the instrument undergoes significant motion (e.g., moving across the image within a few frames). While the MLP decoder loses track of the object during such maneuvers, the Slot-Mixer decoder successfully maintains object tracking. This is evident in the image sequence, where the Slot-Mixer preserves the correct mask order despite the large motion, showcasing its robustness in maintaining temporal coherence.
F.2 Zero-shot Segmentation of longer video sequence
As discussed in Section 4.7 of the main text, the model trained on MICCAI abdominal data demonstrates strong performance on thoracic data, which involves similar instruments but different tissue backgrounds. Figure S4 presents qualitative segmentation results for the zero-shot transfer of Slot-BERT with the Mixer decoder applied to thoracic robotic surgery videos.
Despite the differences in tissue backgrounds, the model effectively segments and tracks instruments. Notably, a slot successfully attends to gauze, an object not present in the MICCAI dataset, as seen in the first, second, and sixth rows. For instruments that are common between the MICCAI abdominal surgery data and the thoracic videos, such as needle drivers, the model achieves superior segmentation results, showcasing its ability to generalize across different surgical contexts.